OpenStudy (anonymous):

Lim (x approaches to 0) of (e^2x -e^x)/sinx Don't understand how to solve this. Don't know how to get 3 (the answer) from this, please help.

OpenStudy (anonymous):

apply l hospitals rule...as it is in 0/0 form

OpenStudy (anonymous):

differentiat the numerator first and then denominator until 0/0 form vanishes and gets reduced to some number.

OpenStudy (anonymous):

have u undrstood?if njot,tell me.i will explain it to u step by step.

OpenStudy (anonymous):

Can you please explain it?

OpenStudy (anonymous):

ok first of all tell me do u know differentiation?

OpenStudy (anonymous):

pls rply fast

OpenStudy (anonymous):

I tried dividing the denominator by x in order to get rid of sin and I got an x for the denominator. But I don't know what to do from there on.

OpenStudy (anonymous):

no no u dont need to.here u have to apply l hodpitals rule.....ok first differentiate the numerator and tell me wt u got.

OpenStudy (anonymous):

Do you mean this? Lim(x->0) ( e^2x )/sinx - (e^-x) /sinx

OpenStudy (anonymous):

in numerator we have e^2x-e^x by differentiating,we get 2e^2x-e^x

OpenStudy (anonymous):

have u undrstood dis diffrentiation?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

in denominator we have sinx.now we have to differentiate it as well, so we get cosx. now, [2e^2x-e^x]/cosx

OpenStudy (anonymous):

now put the value of limit,i.e.x=o in it u will get the ans.

OpenStudy (anonymous):

differentiation of e^ax=a*e^ax do u know dis?

OpenStudy (anonymous):

Yes

OpenStudy (badhi):

If you don't know the L'Hospitals rule try this, \[\lim \limits_{x\to0} \frac{e^{2x}-e^x}{\sin x}=\lim \limits_{x\to0} \frac{e^x(e^{x}-1)}{\sin x}\\=\lim \limits_{x\to0} e^x\frac{(e^{x}-1)}{x}\frac{1}{\left(\frac{\sin x}{x}\right)}\\=\left(\lim \limits_{x\to0}e^x\right)\left( \lim \limits_{x\to0}\frac{(e^{x}-1)}{x}\right)\frac{1}{\left(\lim \limits_{x\to0}\frac{\sin x}{x}\right)}\]

OpenStudy (anonymous):

Which is 1* lim(x->0) e^x -1)/x But where do I go from this?Thr denominator can't be zero

OpenStudy (badhi):

Remember: \[e^x=1+\sum\limits_{i=1}^\infty \frac{x^n}{n!}\\ \frac{e^x-1}{x}=\sum\limits_{n=1}^\infty \frac{x^{n-1}}{n!}=1+\sum\limits_{n=2}^\infty\frac{x^{n-1}}{n!}\\ \lim\limits_{x\to0}\frac{e^x-1}{x}=1+\lim\limits_{x\to0} \sum\limits_{n=2}^\infty\frac{x^{n-1}}{n!}=1+0=1\]

OpenStudy (anonymous):

Ok thanks :)

OpenStudy (anonymous):

l hospital is easier.if u learn abt l hospital u will know it dat its a far better process.

OpenStudy (anonymous):

@Mathstudent97 madrockz is right ...when you solve the complex question then l hospital rule is very easy and time saving compare to another way..if you dont know differentiation then learn it and after it learn l hospital rule then you will observe that l hospital rule is very easy for 0/0 form ..and @madrockz thanks....

OpenStudy (anonymous):

Thanks I will :)